OpenStudy (anonymous):

the homecoming committee bought only 500 plastic souvenir footballs to sell at the homecoming game to raise money for a local charity. The profit (in dollars), p, from the sale of s footballs can be represented by the following equation. p=5s-128 1. Since the homecoming committee bought only 500 footballs, they can sell no more than this. Represent this constraint with an inequality. 2. Is it possible for a profit to be exactly $1500? 3. Is it possible for a profit to be at least $2400?

OpenStudy (thefurball):

Plug in 500 for S. Profit must be equal to or less than that answer.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

So, what is the answer for all three sections?

OpenStudy (thefurball):

you fill in where i put "?"\[p = 5 * 500-128 = ?\]\[p \le ?\] is ? bigger than 1500? 2400? (You find ?)